In a large population, 63 % of the people have been vaccinated. If 4 people are randomly selected, what is the probability that AT LEAST ONE of them has been vaccinated

Algebra ->  Probability-and-statistics -> SOLUTION: Question 1: In a large population, 63 % of the people have been vaccinated. If 4 people are randomly selected, what is the probability that AT LEAST ONE of them has been vaccinated      Log On


   



Question 468803: Question 1: In a large population, 63 % of the people have been vaccinated. If 4 people are randomly selected, what is the probability that AT LEAST ONE of them has been vaccinated?


Answer by ccs2011(207)

The situation where at least one out of 4 are vaccinated includes all but one possibility, that which none of the 4 are vaccinated.
The sum of all probabilities must equal 1.
Let Q be probability none of the 4 are vaccinated.
Let P be probability that at least 1 is vaccinated.
P = 1 - Q
Now if 63% are vaccinated, then 37% are not vaccinated.
Thus any given person has a probability of 0.37 of not being vaccinated.
Q = (0.37)^4 = 0.0187
P = 1 - 0.0187 = 0.981